Nova Company’s total overhead cost at various levels of activity are presented below:

Month Machine-
Hours
Total
Overhead
Cost
April 55,000 $ 183,500
May 45,000 $ 160,500
June 65,000 $ 206,500
July 75,000 $ 229,500

Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 45,000 machine-hour level of activity is:

Utilities (variable) $ 54,000
Supervisory salaries (fixed) 40,000
Maintenance (mixed) 66,500
Total overhead cost $ 160,500

Nova Company’s management wants to break down the maintenance cost into its variable and fixed cost elements.

Required:

1. Estimate how much of the $229,500 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $229,500 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.)

2. Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX.

3. Express the company’s total overhead cost in the form Y = a + bX.

4. What total overhead cost would you expect to be incurred at an activity level of 50,000 machine-hours?

  • Required 1

Estimate how much of the $229,500 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $229,500 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.) (Do not round intermediate calculations.)

  • Required 2

Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX. (Do not round your intermediate calculations. Round the "Variable cost per unit" to 2 decimal places.)

Y = _____ + _____ X

  • Required 3

Express the company’s total overhead cost in the form Y = a + bX. (Do not round your intermediate calculations. Round the "Variable cost per unit" to 2 decimal places.)

Variable Cost per Machine Hour Fixed Cost

Utilities Cost

Supervisory Salaries Cost

Maintenance Cost

Total Overhead Cost

Y=_____ + ______ X

  • Required 4

What total overhead cost would you expect to be incurred at an activity level of 50,000 machine-hours? (Do not round intermediate calculations.)

Total Overhead Cost =
Answer #1
1
Total overhead cost 229500
Less: Utilities 90000 =54000/45000*75000
Less: Supervisory salaries 40000
Maintenance cost in July 99500
2
Machine hours Maintenance cost
High level 75000 99500
Low level 45000 66500
Change 30000 33000
Variable cost 1.10 =33000/30000
Fixed cost 17000 =99500-(75000*1.1)
Y = 17000 + 1.10 X
3
Variable cost per
machine hour
Fixed cost
Utilities cost 1.20
Supervisory salaries cost 40000
Maintenance cost 1.10 17000
Total overhead cost 2.30 57000
Y = 57000 + 2.30 X
4
Total overhead cost 172000 =57000+(50000*2.30)